My daughter is having a tough time understanding math and since I haven't done algebra in years.. I don't know how to help her. She has to solve this word problem using linear systems. Please help.
You and a friend go to a mexican restaurant. You order 2 tacos and 3 enchiladas, and your friend orders 3 tacos and 6 enchiladas. One bill is $8.00 and the other bill is $15.15. How much was each taco and each enchilada?

Answer by askmemath(368)

Let the cost of a taco = $t
and the cost of an enchiladas = $e
Now the cost of 2tacos and 3 enchiladas = 2t+3e = $8 (1)
The cost of 3 tacos and 6 enchiladas = 3t+6e = $15.15 (2)
Multiply (1) with 3 and (2) with 2
6t+9e = 24 (3)
6t+12e = 30.30 (4)
Subtract (3) from (4)
6t+12e-(6t+9e) = 30.30 -24
3e = 6.30
Dividing by 3 throughout
e = $2.10
Substituting in any equation we get t = $0.85
Do let me know if I wasnt clear enough anywhere